# Statistical reduced order modelling for the parametric Helmholtz equation

### v1

@@ 5.2.1 Forward Problem
$\bar{f}(\bm{x},\omega)=\mathrm{exp}(\mathrm{i}kx)=\mathrm{cos}(kx)+\mathrm{i}\;\mathrm{sin}(kx).$ (71)
The covariance in both real and imaginary part is modeled with a Matérn kernel (23) with hyperparameters $\nu=1.5$, $\sigma_{f}=0.8$ and $l_{f}=0.6$, hence a GP is defined for both real and imaginary parts of the forcing. There holds
$s_{\mathrm{re}}$ $\sim\mathcal{GP}(0,k_{f}(\bm{x}))$ (72)
$s_{\mathrm{im}}$ $\sim\mathcal{GP}(0,k_{f}(\bm{x}))$ (73)
@@
$f=\bar{f}+(s_{\mathrm{re}}+\mathrm{i}s_{\mathrm{im}})=\bar{f}+s.$ (75)
Figure 9 shows both real and imaginary parts of the mean incident wave.
Figure 9: Incident wave as forcing term for the 2D scattering example. Real part left, imaginary part right.
@@
It is visible that both the pressure field and its variance are zero at the Dirichlet boundary of the circular scatterer.
For the ROM prior, a moment matching approach with $m=12$ moments matched is chosen. The corresponding error to the FEM full order solution is visible in Figure 11. Using the adjoint error indicator, the ROM error is estimated at the training points which are spread evenly throughout the domain. For that, the full order model does not have to be evaluated in the online phase but only the corresponding adjoint ROM, which is very cheap in comparison. Nevertheless, with the chosen moment matching method, a projection basis has to be built for each training point in the offline phase.
@@
Figure 11: The estimate of the ROM error converges with an increasing number of training points. The quality of the estimation, i.e. the difference of the converged estimator to the reference, depends on the quality of the ROM, cf. Fig. 5-6. In this particular case with $m=12$, an approximation error in the estimator is visible. Increasing $m$ would lead to a better approximation but is linked to higher computational cost.
In this example, the ground truth is not only a more finely discretized version of the FEM prior but involves a model error. The ground truth is found by solving the problem defined with the source term
@@
Figure 12: Data is generated on a more finely discretized mesh ($1131$ degrees of freedom vs. $854$ for the prior). A parameter sample serves as the basis and is superimposed with a sine wave in orthogonal direction to the incoming wave in the prior forcing.
In Table 1, the posterior error is listed for two three different settings: A small data example with $5$ sensors and $20$ observations per sensor, an example with $30$ sensors and $50$ observations and an example with a high number of sensors and observations, i.e. $n_{s}=80$ and $n_{o}=200$. It is visible that in the small data example the remaining error is lowest for the full order problem and highest for the classical approach. The proposed approach with the error estimator yields better results as the classical one. When the number of sensors and observations is increased, the data gain more weight in the Bayesian update compared to the prior and the errors are smaller and more similar throughout the different methods. For the estimated model error standard deviation, the results are very similar throughout the methods and it can’t be clearly told that it is the highest for the posterior without an error estimator.
Table 1: $m=12$, relative posterior error for the different methods for different sensor data configurations. The error for the proposed method with an error estimator yields smaller errors compared to the classical method with only the ROM prior, when there is rather few available data. Increasing the number of sensors and observations decreases the weight of the prior and hence a smaller error is achieved for all methods. The difference in the error between the methods is smaller.
$360\mathrm{Hz}$, $5$ sens, $20$ obs $360\mathrm{Hz}$, $30$ sens, $50$ obs $360\mathrm{Hz}$, $80$ sens, $200$ obs
@@
FOM $1.01\mathrm{e}{-2}$ $8.23\mathrm{e}{-3}$ $8.74\mathrm{e}{-3}$ $7.10\mathrm{e}{-3}$ $8.44\mathrm{e}{-3}$ $6.19\mathrm{e}{-3}$
w/o est $1.39\mathrm{e}{-2}$ $9.17\mathrm{e}{-3}$ $9.78\mathrm{e}{-3}$ $7.64\mathrm{e}{-3}$ $8.48\mathrm{e}{-3}$ $6.30\mathrm{e}{-3}$
With est $1.12\mathrm{e}{-2}$ $8.54\mathrm{e}{-3}$ $8.72\mathrm{e}{-3}$ $7.30\mathrm{e}{-3}$ $8.51\mathrm{e}{-3}$ $6.28\mathrm{e}{-3}$
@@
FOM $1.20\mathrm{e}{-2}$ $1.41\mathrm{e}{-2}$ $4.86\mathrm{e}{-3}$ $1.22\mathrm{e}{-2}$ $1.16\mathrm{e}{-3}$ $1.16\mathrm{e}{-2}$
w/o est $2.85\mathrm{e}{-2}$ $1.54\mathrm{e}{-2}$ $1.36\mathrm{e}{-2}$ $1.20\mathrm{e}{-2}$ $2.10\mathrm{e}{-2}$ $1.21\mathrm{e}{-2}$
With est $1.42\mathrm{e}{-2}$ $1.42\mathrm{e}{-2}$ $4.96\mathrm{e}{-3}$ $1.10\mathrm{e}{-2}$ $1.03\mathrm{e}{-3}$ $9.54\mathrm{e}{-3}$
@@
In Table 2, another setting was chosen: A ROM with more moments matched ($m=20$) and a frequency of $300$Hz, i.e. closer to the expansion frequency of $250$Hz. In this case, the ROM is already a good representation of the underlying full order model and the different methods yield very similar errors. This means that, although with the good ROM the ROM error is estimated accurately, there is no advantage of using the estimator in the inference because the error of the ROM prior is already very small.
Table 2: $m=20$, relative posterior error for the different methods for different sensor data configurations. Because the ROM is already a very good representation of the FOM in every considered parameterization, the error estimator brings no real advantage and the posterior error is similar for all compared methods. Also, very similar results for the estimated model error are found.
$300\mathrm{Hz}$, $5$ sens, $20$ obs $300\mathrm{Hz}$, $30$ sens, $50$ obs $300\mathrm{Hz}$, $80$ sens, $200$ obs
@@
FOM $8.71\mathrm{e}{-3}$ $1.17\mathrm{e}{-2}$ $6.39\mathrm{e}{-3}$ $7.47\mathrm{e}{-3}$ $6.17\mathrm{e}{-3}$ $7.15\mathrm{e}{-3}$
w/o est $8.74\mathrm{e}{-3}$ $1.26\mathrm{e}{-2}$ $6.40\mathrm{e}{-3}$ $7.48\mathrm{e}{-3}$ $6.17\mathrm{e}{-3}$ $7.15\mathrm{e}{-3}$
With est $8.74\mathrm{e}{-3}$ $1.18\mathrm{e}{-2}$ $6.40\mathrm{e}{-3}$ $7.49\mathrm{e}{-3}$ $6.17\mathrm{e}{-3}$ $7.15\mathrm{e}{-3}$
@@
FOM $3.16\mathrm{e}{-4}$ $9.19\mathrm{e}{-3}$ $4.09\mathrm{e}{-3}$ $1.22\mathrm{e}{-2}$ $1.12\mathrm{e}{-3}$ $1.15\mathrm{e}{-2}$
w/o est $3.09\mathrm{e}{-4}$ $6.34\mathrm{e}{-3}$ $4.05\mathrm{e}{-3}$ $1.21\mathrm{e}{-2}$ $1.12\mathrm{e}{-2}$ $1.15\mathrm{e}{-2}$
With est $3.09\mathrm{e}{-4}$ $8.84\mathrm{e}{-3}$ $3.97\mathrm{e}{-3}$ $1.22\mathrm{e}{-2}$ $1.12\mathrm{e}{-2}$ $1.15\mathrm{e}{-2}$
@@ 6 Conclusion
A ROM adaptation of statFEM has been developed. The Bayesian prior is not constructed using FEM but with a ROM based on a moment matching approach. This is necessary if the frequency dependent full order model is too expensive to be solved at many frequency steps. Compared to the classical approach applied to a ROM prior, statROM performs better because the ROM error is estimated beforehand. This was shown by computing the error and convergence properties of the classical and the proposed approach compared to a full order reference solution. The new method enables solving the prior in frequency sweeps much faster as it is the case for a full order model while estimating the ROM error.
@@
As a next step, efforts are directed towards parameterizing the ROM, which currently can be applied primarily to frequency sweeps. Another important extension to the method is finding a way to keep the memory requirements low, especially for the covariance matrices in larger systems. As soon as parallelization on the frequency grid is applied, several large matrices have to be kept in memory. E.g., the approach by [41], where sparse precision matrices replace their dense covariance counterpart, could be adapted to work in the setting presented here. Also, a practical application in aerospace acoustics is considered as a feasible applied problem for statROM.
@@
    All code was implemented in Python using FEniCSx [42, 43, 44, 45] as FEM backend. The source code to generate the results in this publication is available at https://github.com/herluc/statROM [46].

### v2

@@ 5.2.1 Forward Problem
$\bar{f}(\omega,\bm{x})=\mathrm{exp}(\mathrm{i}kx)=\mathrm{cos}(kx)+\mathrm{i}\;\mathrm{sin}(kx).$ (74)
The mean of the incident wave is constant along the y axis. The covariance in both real and imaginary part is modelled with a Matérn kernel (22) with hyperparameters $\nu=1.5$, $\sigma_{f}=0.8$ and $l_{f}=0.6$, hence a GP is defined for both real and imaginary parts of the forcing. Therefore, we have
$f_{GP_{\mathrm{re}}}$ $\sim\mathcal{GP}(0,k_{f}(\bm{x}))$ (75)
$f_{GP_{\mathrm{im}}}$ $\sim\mathcal{GP}(0,k_{f}(\bm{x}))$ (76)
@@
$f=\bar{f}+(f_{GP_{\mathrm{re}}}+\mathrm{i}f_{GP_{\mathrm{im}}})=\bar{f}+f_{GP}.$ (77)
The stochastic forcing and subsequent sampling lead to an incident wave that is not constant along the y axis any more, compared to the mean.
@@
One can observe that both the pressure field and its variance are zero at the Dirichlet boundary of the circular scatterer. The prior computation is the step in the workflow where using the ROM is crucial: The solver time per sample is significantly reduced compared to the full order model. For the 2D example, we measured an average of $0.001938\text{s}$ per sample for the full order model and an average of $0.0006217\text{s}$ per sample for the ROM. This includes reduction of the matrices, solving the reduced system and projecting back onto the full order space. FE system assembly has to be done for both the full order model and the ROM and was excluded from both measurements. Per training point in the estimator, one solution of the reduced order system is performed. Therefore, one has to assess whether the statROM approach is still faster than the full order approach when many error estimator training points are in use. In the case of the 2D example with $256$ parameter samples, the total prior runtime for the full order approach amounts to $6.4382\text{s}$ and for the statROM approach $5.8377\text{s}$, including the error estimator with $200$ training points ($0.29373\text{s}$) and FE assembly. Hence, in this example, $256$ FEM solves are compared to $456$ ROM solves. Note that the computation of the error estimator could be accelerated further by e.g. using sparse GP approximation [51] and the global Arnoldi method, see Remark 2.
For the ROM prior, a moment matching approach with $m=12$ is chosen. The corresponding error to the FEM full order solution is visible in Figure 9. Using the adjoint error indicator, the ROM error is estimated at the training points which are spread evenly throughout the domain. For that, the full order model does not have to be evaluated in the online phase but only the corresponding adjoint ROM, which is cheaper in comparison.
@@
Figure 9: The estimate of the ROM error improves with an increasing number of training points. The quality of the estimation, i.e. the difference of the converged estimator to the reference, depends on the quality of the ROM, cf. Fig. 5. In this particular case with $m=12$, an approximation error in the estimator is visible. Increasing $m$ would lead to a better approximation but is linked to higher computational cost.
In this example, the ground truth is not only a more finely discretized version of the FEM prior but involves a different modelling choice as well. In particular, the ground truth is found by solving the problem defined with the source term
@@
Figure 10: Data are generated on a more finely discretized mesh ($1131$ degrees of freedom vs. $854$ for the prior). A parameter sample (left) serves as the basis and is superimposed with a sine wave in orthogonal direction to the incoming wave in the prior forcing (right)
In Table 1, the posterior error is listed for three different sensor configurations: low resolution, moderate resolution, and high resolution. The low resolution configuration is defined with $5$ sensors and $20$ observations per sensor, the moderate resolution one with $30$ sensors and $50$ observations and the high resolution configuration with $n_{s}=80$ and $n_{o}=200$.
Table 1: $m=12$, relative posterior error norm for the different methods for different sensor data configurations. The respective posterior mean is compared to the ground truth. The proposed method with an error estimator yields smaller errors compared to the standard method with the ROM prior, if the size of the data set is small. Increasing the number of sensors and observations decreases the weight of the prior and hence a smaller error is achieved for all methods. The difference in the error between the methods is smaller.
$400\mathrm{Hz}$, $5$ sens, $20$ obs $400\mathrm{Hz}$, $30$ sens, $50$ obs $400\mathrm{Hz}$, $80$ sens, $200$ obs
@@
FOM $2.89\mathrm{e}{-2}$ $1.77\mathrm{e}{-2}$ $1.03\mathrm{e}{-2}$ $8.75\mathrm{e}{-3}$ $1.03\mathrm{e}{-2}$ $7.70\mathrm{e}{-3}$
w/o est $5.76\mathrm{e}{-2}$ $4.34\mathrm{e}{-2}$ $1.14\mathrm{e}{-2}$ $06\mathrm{e}{-2} 1.1$$10\mathrm{e}{-2}$ $7.98\mathrm{e}{-3}$
With est $3.19\mathrm{e}{-2}$ $2.63\mathrm{e}{-2}$ $1.04\mathrm{e}{-2}$ $1.05\mathrm{e}{-2}$ $1.07\mathrm{e}{-2}$ $8.02\mathrm{e}{-3}$
@@
FOM $2.17\mathrm{e}{-2}$ $3.44\mathrm{e}{-2}$ $1.93\mathrm{e}{-3}$ $7.15\mathrm{e}{-3}$ $2.96\mathrm{e}{-3}$ $9.72\mathrm{e}{-3}$
w/o est $1.96\mathrm{e}{-2}$ $4.52\mathrm{e}{-2}$ $1.60\mathrm{e}{-2}$ $1.25\mathrm{e}{-2}$ $1.78\mathrm{e}{-2}$ $7.99\mathrm{e}{-3}$
With est $2.29\mathrm{e}{-2}$ $3.51\mathrm{e}{-2}$ $1.29\mathrm{e}{-2}$ $6.99\mathrm{e}{-3}$ $3.92\mathrm{e}{-3}$ $5.87\mathrm{e}{-3}$
@@
Note that in the low resolution example the remaining error is lowest for the full order problem and highest for the standard statFEM on ROM prior approach, i.e. without the error estimator. The proposed approach with the error estimator yields better results as the standard one. When the number of sensors and observations is increased, the data gain more weight in the Bayesian update compared to the prior and the errors are smaller and more similar throughout the different methods. For the estimated model error standard deviation, the results are very similar throughout the methods. For the low resolution configuration, the differences of the respective posterior state to the ground truth, i.e. reference solution, are plotted in Figure 11. Notice that the difference is the largest for the standard statFEM performed on the ROM prior. The proposed approach yields a difference closer to the full order approach.
Figure 11: The difference of the ground truth (reference) mean $\mu_{u\text{;ref}}$ to the posterior mean of the full order method (left column), the standard statFEM approach applied to a ROM prior (middle column) and the proposed statROM method (right column). We observe that the posterior mean of the proposed reduced order method ($\mu_{\tilde{u}|{\bm{Y};d_{r}}}$) yields results closer to the full order statFEM method ($\mu_{{u}|\bm{Y}}$) compared to the standard statFEM on a ROM prior ($\mu_{\tilde{u}|{\bm{Y}}}$). This holds true both for the real (top row) and imaginary (bottom row) parts of the solution.
In Table 2, the posterior error is listed with the full order posterior as reference. As before, we observe that the proposed approach yields a lower error compared to the standard approach. The difference plots for the low resolution configuration in Figure 12 support this observation.
Table 2: $m=12$, relative posterior error norm for the different methods for different sensor data configurations. The respective posterior mean is compared to the full order counterpart. The proposed method with an error estimator yields smaller errors compared to the statFEM method with a ROM prior, when the data size is rather small. By increasing the number of sensors and observations, the weight of the prior is decreased and hence, a smaller error is achieved for all methods. Also the difference in the error between the methods is smaller.
$400\mathrm{Hz}$, $5$ sens, $20$ obs $400\mathrm{Hz}$, $30$ sens, $50$ obs $400\mathrm{Hz}$, $80$ sens, $200$ obs
@@
w/o est $1.23\mathrm{e}{-2}$ $03\mathrm{e}{-2} 1.1$$18\mathrm{e}{-3}$ $2.11\mathrm{e}{-3}$ $5.14\mathrm{e}{-4}$ $5.49\mathrm{e}{-4}$
With est $1.73\mathrm{e}{-3}$ $3.27\mathrm{e}{-3}$ $5.37\mathrm{e}{-4}$ $1.22\mathrm{e}{-3}$ $3.87\mathrm{e}{-4}$ $6.88\mathrm{e}{-4}$
@@
Figure 12: The posterior mean difference in the low resolution sensor configuration of the full order solution and the standard statFEM approach (top row) compared to the proposed statROM approach (bottom row). We observe that the posterior mean of the proposed reduced order method ($\mu_{\tilde{u}|{\bm{Y};d_{r}}}$) yields results much closer to the full order statFEM method ($\mu_{{u}|\bm{Y}}$) than the standard statFEM ($\mu_{\tilde{u}|{\bm{Y}}}$) does. This holds true both for the real and imaginary parts of the solution.
In Figure 13, a slice through the 2D posterior states is plotted so that the difference between the proposed method, the standard method and the full order reference can be easily observed. Although the errors are small compared to the 1D example, notice that the mean of the proposed statROM method approaches the mean of the full order method, whereas the error for the standard approach is higher. The variance is similar throughout the different methods.
Figure 13: A slice in x-direction through the posterior states at $y=0.75$. Top: Real part, Bottom: Imaginary part. The full order reference is plotted in blue, the standard statFEM method applied to a ROM prior is plotted in yellow, and the proposed approach is plotted in purple. We observe that the mean of the proposed statROM approach comes closer to the full order posterior than the standard approach does. Notice that the standard deviation is similar for all methods.
Choosing another setting, i.e. a ROM with more moments matched ($m=20$) and a frequency of 300300Hz, i.e. closer to the expansion frequency of 250250Hz, the ROM already provides a good representation of the underlying full order model and the different methods yield very similar errors. This means that, although the ROM error is estimated accurately, there is no advantage of using the estimator in the inference because the error of the ROM prior is already very small. We omit a table for this example.
In Table 3, the posterior error is listed with the full order posterior as reference for the $300\mathrm{Hz}$, $m=20$ setting where we also observe no large differences between the methods.
Table 3: $m=20$, relative posterior error norm for the different methods for different sensor data configurations. The respective posterior mean is compared to the full order counterpart. The proposed method with an error estimator yields smaller errors compared to the standard statFEM method with the ROM prior, in case of a small data set. Increasing the number of sensors and observations decreases the weight of the prior and hence a slightly smaller error is achieved for all methods.
$300\mathrm{Hz}$, $5$ sens, $20$ obs $0\mathrm{Hz}, 3030$ sens, $50$ obs $300\mathrm{Hz}$, $80$ sens, $200$ obs
Method err re err im err re err im err re err im
w/o est $3.02\mathrm{e}{-6}$ $1.43\mathrm{e}{-5}$ $2.53\mathrm{e}{-6}$ $1.55\mathrm{e}{-5}$ $1.60\mathrm{e}{-6}$ $5.79\mathrm{e}{-6}$
With est $1.42\mathrm{e}{-6}$ $2.33\mathrm{e}{-6}$ $2.84\mathrm{e}{-6}$ $3.75\mathrm{e}{-6}$ $8.90\mathrm{e}{-7}$ $2.33\mathrm{e}{-6}$
@@
We want to summarize the findings from the 2D example and point out the most important building blocks of the method as well as the approximations that had to be made:
  1. 1.
    QMC prior construction
    • The QMC method approximates the solution statistics of the given problem. Approximation errors occur because of a limited sample size and the Gaussian assumption.
    • For all QMC samples, a ROM basis was constructed which serves as a fast-to-solve approximation to the full order counterpart.
  2. 2.
    ROM error estimation
    • The ROM approximation error was estimated using the adjoint approach. An adjoint ROM basis was constructed in the offline phase.
    • Errors occur because of a limited adjoint ROM basis size and a limited amount of training points for the GP regression.
  3. 3.
    Data Assimilation
    • Here we compared three different settings. 1. The full order statFEM, 2. statFEM using the ROM prior without an error estimator and 3. The proposed statROM approach with an error estimator.
    • All settings suffer from noisy data. Also, with limited data and a mismatch between prior and data, there is still an error between the posterior and the ground truth.
    • In the chosen settings, the ROM prior exhibits a stronger mismatch to the data, which is why the posterior error is higher than for the full order method.
    • The key finding of this paper is that the proposed statROM method is able to mitigate that higher error to some extend. This makes it possible to solve the QMC ensemble more quickly using the ROM, while maintaining a high accuracy in the posterior.
@@ 6 Conclusion
A ROM adaptation of statFEM has been developed. The Bayesian prior is not constructed using FEM but with a ROM based on a moment matching approach. This is necessary if the frequency dependent full order model is too expensive to be solved at many frequency steps. Compared to the standard statFEM approach applied to a ROM prior, statROM performs better because the ROM error is estimated beforehand. This was shown by computing the error and convergence properties of the standard and the proposed approach compared to a full order reference solution. The new method enables solving the prior in frequency sweeps much faster as it is the case for a full order model while estimating the ROM error.
@@
As a next step, efforts are directed towards parameterizing the ROM, which currently can be applied primarily to frequency sweeps. Another important extension to the method is finding a way to keep the memory requirements low, especially for the covariance matrices in larger systems. As soon as parallelization on the frequency grid is applied, several large matrices have to be kept in memory. For instance, the approach by [52], where sparse precision matrices replace their dense covariance counterpart, could be adapted to work in the setting presented here. Also, a practical application in aerospace acoustics is considered as a feasible applied problem for statROM.
@@
    All code was implemented in Python using FEniCSx [53, 54, 55, 56] as FEM backend. The source code to generate the results in this publication is available at https://github.com/herluc/statROM and on Zenodo https://doi.org/10.5281/zenodo.15746081 [57]. The code is able to reproduce both the 1D and the 2D scattering example. It returns information similar to what is presented in Tables 1, 2 and 3 as well as a selection of plots. The user can set the most important parameters and observe the change in the posterior solution quality. One limitation of the code lies in its ability to generalize to other examples: Right now, only the given Helmholtz PDE can be used on the given geometries, so it would take significant effort to switch to another PDE. Also, the convergence studies are not automated within the code.